OpenStudy (anonymous):

Simplify square root of 3 multiplied by the fifth root of 3

OpenStudy (anonymous):

3 to the power of 1 over 10 3 to the power of 3 over 5 3 to the power of 9 over 10 3 to the power of 7 over 10

OpenStudy (campbell_st):

write them in index from \[3^{\frac{1}{2}} \times 3^{\frac{1}{5}}\] the rule for multiplying the same base is add the powers \[x^a \times x^b = x^{a + b}\] you need to apply the rule to this question.

OpenStudy (acxbox22):

1/2+1/5=2/10+5/10 7/10 \[3^{7/10}\]

OpenStudy (anonymous):

square root of 3 multiplied by the fifth root of 3 is question and the next post is the answers that where giving